OpenStudy (anonymous):

Find the tangent line approximation to 12/x near x = 12?

OpenStudy (tkhunny):

Can you find the 1st derivative of f(x) = 12/x?

OpenStudy (anonymous):

-12/x^2 right?

OpenStudy (tkhunny):

That's it. f(x) = 12/x f'(x) = -12/x^2 Now you need to evaluate those at x = 12 and you will have all the information you need.

OpenStudy (anonymous):

So Im I trying to find f(12)+fprime(12)(x-12)?

OpenStudy (tkhunny):

Yes.

OpenStudy (anonymous):

But I dont get where do i get the x from?

OpenStudy (tkhunny):

Problem statement? x = 12

OpenStudy (anonymous):

Well the equation we got in class is f(a)+fprime(a)(x-a)

OpenStudy (anonymous):

and the teacher said that a is equal to the x your are given

OpenStudy (anonymous):

so i get 1+(-.083(x-12)

OpenStudy (tkhunny):

1) Why would you convert it to decimals? 2) Who cares about 'a' or 'fred" or whatever. You need the equation of a line. You have a point: (12,f(12)) = (12,1) You have a slope: f'(12) = -1/12 That's all you need for a line if you remember the Point-Slope Form. There is just no need to memorize yet another form for "linearization". You already know how to write the equation of a line.

OpenStudy (anonymous):

So it should be y=13/2(X-12)?

OpenStudy (tkhunny):

Well, let's see... \(y - 1 = -\dfrac{1}{12}(x-12)\) \(y = -\dfrac{1}{12}(x-12) + 1\) Why do we have to write more than that? \(12y = -(x-12) + 12\) \(12y = -x + 12 + 12\) \(12y = -x + 24\) \(x + 12y = 24\) \(\dfrac{x}{24} + \dfrac{y}{2} = 1\) Lots of ways to write it. Which do you find suitable?

OpenStudy (anonymous):

okay yeah thanks it was just showing up wrong because I wasn't supposed to enter it as y=

OpenStudy (tkhunny):

Perfect.